# Dark Photon Dark Matter and Low-Frequency Gravitational Wave Detection with Gaia-like Astrometry

### v1

@@ II Background motion of the apparent stellar locations
Our result is shown in Fig. 1, where one can observe that the linear subtraction approach effectively reduces the noise from stellar proper motion, reproducing the PSD of pure white noise (WN) above $10^{-8}$ Hz. However, for lower frequency, part of the WN is also removed by the linear subtraction, leading to a smaller PSD. Furthermore, the inclusion of binaries indeed introduces significant noise when $f\lesssim 10^{-7}\,{\rm Hz}$, corresponding to $P_{b}\gtrsim 24$ day. This noise grows as $P_{b}$ increases, coinciding with the behavior manifested by Eq. (5). As $f$ falls below $6\times 10^{-9}\,{\rm Hz}$, the observed angular deflection $A_{b}$ of the binary system no longer exhibits periodic oscillations during the duration of Gaia’s observations. Consequently, two subtraction schemes will reduce the time-dependent noise induced by unresolved binaries in the low-frequency region, with the quadratic subtraction being more efficient. However, the quadratic subtraction also leads to more reductions to our signal at low frequencies, and we will see that the linear subtraction achieves better signal-to-noise ratio eventually.
III Astrometric deflections induced by GW and DPDM
III.1 GWs from binary systems
To begin with, we first revisit the formula for light deflection induced by GWs Book and Flanagan (2011); Pyne et al. (1996). The presence of GWs triggers metric perturbations, which lead to slight deflection of light traveling from the star to the Earth, ultimately resulting in the apparent displacement of stellar positions. By repeatedly measuring the apparent positions of a large number of stars across the sky, Gaia might be capable of capturing such deflection and thus extracting the characteristic patterns of GWs.
The variation in the line of sight direction $n^{i}$ due to GWs can be expressed as Pyne et al. (1996); Book and Flanagan (2011)
$\delta n^{i}\simeq\frac{n^{i}-p^{i}}{2(1-\bm{n}\cdot\bm{p})}h_{jk}({\rm E})n^{j}n^{k}-\frac{1}{2}h_{ij}({\rm E})n^{j}\;,$ (8)
where the unit vector $\bm{p}$ described by two rotational angles $\theta_{p}$ and $\varphi_{p}$ represents the propagated direction with $p^{i}$ being its components, and $h_{ij}(\rm E)$ refers to the metric perturbation calculated at the Earth. The metric perturbation $h_{ij}$ caused by monochromatic GWs with frequency $f_{\rm GW}$ reads
$h_{ij}(t,\boldsymbol{x})={\rm Re}\left[\mathcal{H}^{\rm GW}_{ij}e^{2\pi if_{\rm GW}(t-{\bm{p}}\cdot\bm{x})}\right]\;,$ (9)
@@ III.1 GWs from binary systems
$\mathcal{H}^{\rm GW}_{ij}=h_{+}\epsilon^{+}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{+}}+h_{\times}\epsilon^{\times}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{\times}}\;,$ (10)
where $\epsilon^{+}_{ij}$ and $\epsilon^{\times}_{ij}$ are two polarization tensors depending on $\boldsymbol{p}$ and the GW polarization angle $\psi$ Arzoumanian et al. (2023), $h_{+}$ and $h_{\times}$ are their amplitudes, and $\phi_{+}$ and $\phi_{\times}$ are two corresponding phases. We consider the monochromatic GW radiated from a binary system with a circular orbit, the amplitudes $h_{+}$ and $h_{\times}$ of which can be respectively expressed as
$h_{+}=h^{\rm GW}_{0}\frac{1+\cos^{2}\iota}{2}\;,\quad h_{\times}=h^{\rm GW}_{0}\cos\iota\;,$ (11)
where $h^{\rm GW}_{0}$ is a constant depending on the chirp mass, the angular orbital frequency and the luminosity distance to the source, and $\iota$ is the inclination angle of the binary. Finally, for the circular orbit, the relation $\varphi_{+}-\varphi_{\times}=\pi/2$ should be satisfied.
We emphasize that we adopt the distant source approximation, thus we only need to keep the “Earth term” in the metric perturbation in Eq. (8). This approximation is valid when the sources are located at distances much larger than the wavelength of GWs Book and Flanagan (2011); Pyne et al. (1996); Moore et al. (2017). Under this approximation, the perturbations at different sources are uncorrelated, resembling a kind of random noise. Such noise is much smaller than the Gaia’s resolution and can be safely omitted.
III.2 DPDM with purely gravitational interactions
Analogous to GWs, dark photon fields can also induce metric fluctuations via gravitational effects Nelson and Scholtz (2011); Khmelnitsky and Rubakov (2014); Nomura et al. (2020). In this subsection, we derive the angular deflection of light emitted from the star as it passes through DPDM with purely gravitational interactions due to metric perturbations Çalışkan et al. (2024). We should mention that DPDM with a mass below $10^{-22}$ eV is subject to severe constraints from, e.g., the integrated Sachs-Wolfe effect on CMB anisotropies ($\gtrsim 10^{-24}~{}{\rm eV}$Hlozek et al. (2015), the stellar kinematics of dwarf spheroidal galaxies (a few times of $10^{-23}$ eV) González-Morales et al. (2017); Kendall and Easther (2020) and the Lyman-$\alpha$ forest ($\gtrsim 10^{-21}~{}{\rm eV}$Hui et al. (2017); Armengaud et al. (2017). Though the lower bound on ultralight dark matter mass is still under debate, here we assume that dark photons with $m_{A}\lesssim 10^{-22}~{}{\rm eV}$ constitute at most 10% of the total dark matter energy density.
@@ III.2 DPDM with purely gravitational interactions
where the polarization vector $\boldsymbol{q}(\boldsymbol{x})=(\cos\theta_{q}\cos\varphi_{q}e^{i\phi_{1}}$, $\cos\theta_{q}\sin\varphi_{q}e^{i\phi_{2}},\sin\theta_{q}e^{i\phi_{3}})^{\rm T}$ is parametrized by two polarization angles $\theta_{q}$ and $\varphi_{q}$ and three phases $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, all of which are in general certain functions of $\boldsymbol{x}$. Under the synchronous gauge, the metric perturbation has the following form
$h_{ij}^{\rm osc}(t,\boldsymbol{x})\simeq{\rm Re}\left[\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})e^{2im_{A}t}\right],$ (13)
@@
$\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})=-h^{\rm osc}_{0}\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right].$ (14)
The signal amplitude can be written as $h^{\rm osc}_{0}=4\pi GA^{2}_{0}(\boldsymbol{x})$, with $G$ and $A_{0}(\boldsymbol{x})$ being respectively the gravitational constant and the amplitude of dark photon field. The amplitude of DPDM is
@@
$h^{\rm osc}_{0}=5.2\times 10^{-15}\left(\frac{\rho_{\rm DM}}{0.4\,{\rm GeV}/{\rm cm}^{3}}\right)\left(\frac{10^{-23}\,{\rm eV}}{m_{A}}\right)^{2}.$ (16)
It is known that the metric perturbation $h_{ij}$ given in Eq. (13) yields a delay of the arrival time of the pulsar light, which can be probed by PTAs Nomura et al. (2020); Xue et al. (2022); Wu et al. (2022); Xia et al. (2023). Here we show that the same metric perturbation could also lead to the deflection of star’s proper position, namely,
$\delta n^{i}\simeq-\frac{(\delta^{ik}-n^{i}n^{k})}{2}n^{j}h_{jk}({\rm E})+\mathcal{O}(v_{\rm vir})\;,$ (17)
where the perturbation $h_{jk}({\rm E})$ is again calculated at the Earth. One can obtain the above expression by taking the limit $\boldsymbol{p}\sim v_{\rm vir}\ll 1$ in Eq. (8). The complete gauge-invariant expressions of $\delta n^{i}$ can be found in appendix B. In this case the star terms are suppressed by the distance and the virial velocity, as can be seen from Eqs. (71) and (76).
III.3 DPDM with additional gauge couplings
DPDM can also interact with ordinary matter if they both carry ${\rm U(1)}_{B}$ or ${\rm U(1)}_{B-L}$ charge. As a consequence, any object could experience an oscillating force due to the DPDM background Pierce et al. (2018)
@@ III.3 DPDM with additional gauge couplings
which can cause aberrations of the stellar proper motion due to the moving observer Guo et al. (2019). The resulting apparent angular deflection can be expressed as
$\delta\boldsymbol{n}(t,\boldsymbol{n})=\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})-(\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})\cdot\boldsymbol{n})\boldsymbol{n}\;.$ (20)
where $\boldsymbol{x}_{\rm E}$ is the coordinate of the observer. Moreover, the dark photon field can also directly lead to stellar motion, with an amplitude approximated by $\delta\boldsymbol{n}\lesssim|\delta\boldsymbol{v}|\cdot t/D$. However, for the typical observation time $(\sim 1~{}{\rm year})$ and distance $D\gtrsim 1~{}{\rm kpc}$, $\delta\boldsymbol{n}$ is rather tiny and will not be taken into consideration Guo et al. (2019).
IV Sensitivity of GW and DPDM searches with Gaia
In this section, we estimate the sensitivity of Gaia-like astrometry for GW and DPDM induced signals. We consider both Gaia and its next-generation upgrade, referred to as “XG-Gaia”. The operational parameters of these missions are set as follows:
@@ IV Sensitivity of GW and DPDM searches with Gaia
where $\sigma_{\rm ss}$ is the noise from a single exposure, [^2] The uncertainties on the proper motion parameters depend on stellar magnitudes $G$. It is expected the final uncertainty $\sigma_{\rm ss}\simeq 100~{}\mu{\rm as}$ for stars with $G\simeq 20$ in Gaia DR5 Vallenari et al. (2023). As these stars constitute the vast majority of the Gaia dataset, we will use $\sigma_{\rm ss}=100~{}\mu{\rm as}$ for all stars later on. $\Delta t$ is the observational cadence, $T_{\rm obs}$ is the total observation time, and $N$ is the total number of stars considered in the analysis.
@@
As previously mentioned, we compress the data of $10^{5}$ stars in the same sky area by randomly generating $\widetilde{N}=10^{4}$ virtual stars uniformly distributed across the sky. Consequently, the instrument error for the measurement on each virtual star is rescaled to $\widetilde{\sigma}=\sigma_{\rm ss}/\sqrt{10^{9}/10^{4}}$. To estimate the sensitivity in detecting GWs and DPDM with Gaia-like experiments, we construct a mock dataset by injecting a specific GW or DPDM signal into the virtual stars. In all cases, it has been shown that the signals we considered are always dominated by the Earth term. Therefore, the astrometric effects on the stars in one area, caused by GWs or DPDM, can be approximated to the global effect on the virtual stars within that area.
@@
In addition to the signal component, in Sec. II we have established that the background motion due to stellar proper motion and the acceleration of the SSB is effectively removed by linear subtraction. The constant acceleration resulting from the galaxy’s mass distribution is sufficiently small, thus can be safely ignored. Since linear subtraction is consistently applied to the astrometric data of virtual stars, background motion is not injected into the analysis. To address the time-dependent noise caused by stellar binaries, we adopt two different strategies for Gaia and XG-Gaia:
@@
Figure 2: The 95% exclusion limits from Gaia (green lines) and XG-Gaia (red lines) on the GW strain $h_{0}$ from a binary system, the energy density $\rho_{\rm DM}$ of the DPDM with purely gravitational interactions, and the coupling constant $\epsilon$ of $U(1)_{B}/U(1)_{B-L}$ DPDM, respectively. Model parameters regarding all the angles have been integrated out. Solid lines represent the bounds with linear subtraction, while dashed lines correspond to quadratic subtraction. In addition, we use blue lines to denote the results including the time-dependent noise from unresolved binary systems. For comparison, constraints from NANOGrav 15-year Agazie et al. (2023b); Afzal et al. (2023), EPTA DR2 Smarra et al. (2023), and PPTA DR2 Xue et al. (2022); Wu et al. (2022) results are labeled by orange, cyan and purple curves, respectively, with the solid (dashed) curves being the (un)correlated constraints. In the top-right panel, the gray dashed line shows the amount of local dark matter density near the Earth as a reference. In the bottom two panels, the horizontal solid lines are the limits from the MICROSCOPE WEP experiment Touboul et al. (2017, 2022); Amaral et al. (2024).
@@
  • For Gaia, we assume that all stellar binaries are unresolvable. Thus we use the PSD $S_{\rm bin}(f)$ calculated in Sec. II, as shown in Fig. 1.
@@
$\begin{split}\text{(A)}:&\quad\{f_{\rm GW},\;h_{0},\;\iota,\;\phi_{+},\;\theta_{p}\;,\varphi_{p},\;\psi\}\;;\\ \text{(B)}:&\quad\{m_{A},\;\rho_{\rm DM},\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\}\;;\\ \text{(C)}:&\quad\{m_{A},\;\epsilon,\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\}\;.\end{split}$
@@
We fix the values of $f_{\rm GW}$ and $m_{A}$, and scan the other model parameters in their individual parameter space in each case. The 95% exclusion limits on the $h_{0}$, $\rho_{\rm DM}$ and $\epsilon$ after integrating out all angular parameters are shown in Fig. 2. Some remarks are in order. Firstly, in the high-frequency range ($f\gtrsim 2/T_{\rm obs}$), the limits obtained using two different subtraction methods are indistinguishable, as expected, whereas linear subtraction offers better sensitivity in the low-frequency range compared to quadratic subtraction. In particular, the 95% exclusion limits are enhanced by more than one order of magnitude at $f_{\rm GW}\lesssim 6\times 10^{-10}\,{\rm Hz}$ for GWs, or at $m_{A}\lesssim 3\times 10^{-24}\,{\rm eV}$ for DPDM. Secondly, the impact of time-dependent noise from unresolved binary systems significantly decreases the sensitivity by up to an order of magnitude. Hence the inclusion of such noise is important in our analysis. In addition, this decrease is more severe in the linear subtraction scheme than in the quadratic subtraction, as we explained in Sec. II. Lastly, XG-Gaia will possess extended observation duration and improved precision. It will also have enhanced resolution capabilities for binary systems, which is crucial for suppressing the noise from unresolved binaries. It is expected that the sensitivity of XG-Gaia in detecting GWs and DPDM can be increased by more than three orders of magnitude.
V Conclusion
Astrometric measurements offer a new approach to detecting signals such as GWs and DPDM in the nHz frequency range, which are independent of PTA experiments. Subtracting stellar background motion is crucial for accurately estimating the sensitivity of such surveys. In this paper, we show that the background motion of stars can be well described by their proper motion with negligible acceleration, which motivates us to propose a linear subtraction scheme. Compared to the commonly used quadratic subtraction, linear subtraction significantly enhances the detection sensitivity in the low-frequency range. We also find that the unsolved stellar binary systems may produce large time-dependent noise, reducing the detection capability. We thus expect that the next-generation upgrade of Gaia will not only achieve higher precision but also provide better resolution of binary systems in the Milky Way.

### v2

@@ 2 Background motion of the apparent stellar locations
We present the average noise power spectral density (PSD) in Fig. 1, together with error bars of each bin. The uncertainties of PSDs estimated from the simulation are around $10^{-19}/{\rm Hz}^{-1}$. In the absence of the noise from stellar binaries, the linear/quadratic subtraction approach effectively reduces the noise from stellar proper motion, reproducing the PSD of pure white noise above $2/T_{\rm obs}\sim 6\times 10^{-9}$ Hz. Nevertheless, the inclusion of binaries indeed introduces significant noise when $f\lesssim 10^{-7}\,{\rm Hz}$, corresponding to $P_{b}\gtrsim 24$ day. This noise grows as $P_{b}$ increases, coinciding with the behavior manifested by Eq. (5). Furthermore, there are clear breaks in the PSDs at $f=2/T_{\rm obs}$, because the observed angular deflection of isolated stars or binary systems no longer exhibits periodic oscillations during the duration of Gaia’s observations if $f\lesssim 2/T_{\rm obs}$, and consequently, two subtraction schemes will reduce the time-dependent noise in the low-frequency region. The quadratic subtraction removes both the linear and quadratic components of the noise, resulting in lower noise PSDs compared to the linear subtraction. However, the quadratic subtraction also leads to more reductions to our signal at low frequencies, and we will see that the linear subtraction achieves better signal-to-noise ratio eventually.
3 Astrometric deflections induced by GW and DPDM
3.1 GWs from binary systems
To begin with, we revisit the formula for light deflection induced by GWs (Book & Flanagan (2011); Pyne et al. (1996)). The presence of GWs triggers metric perturbations, which lead to slight deflection of light traveling from a star to the Earth, ultimately resulting in the apparent displacement of stellar positions. By repeatedly measuring the apparent positions of a large number of stars across the sky, Gaia might be capable of capturing such deflection and thus extracting the characteristic patterns of GWs.
The variation in the line of sight direction $n^{i}$ due to GWs can be expressed as (Pyne et al. (1996); Book & Flanagan (2011))
$\delta n^{i}_{\rm GW}\simeq\frac{n^{i}-p^{i}}{2(1-\bm{n}\cdot\bm{p})}h^{\rm GW}_{jk}({\rm E})n^{j}n^{k}-\frac{1}{2}h^{\rm GW}_{ij}({\rm E})n^{j}\;,$ (8)
where the unit vector $\bm{p}$ described by two rotational angles $\theta_{p}$ and $\varphi_{p}$ represents the propagated direction with $p^{i}$ being its components, and $h^{\rm GW}_{ij}(\rm E)$ refers to the metric perturbation calculated at the Earth. The metric perturbation $h^{\rm GW}_{ij}$ caused by monochromatic GWs with frequency $f_{\rm GW}$ reads
$h^{\rm GW}_{ij}(t,\boldsymbol{x})={\rm Re}\left[\mathcal{H}^{\rm GW}_{ij}e^{2\pi if_{\rm GW}(t-{\bm{p}}\cdot\bm{x})}\right]\;,$ (9)
@@ 3.1 GWs from binary systems
$\mathcal{H}^{\rm GW}_{ij}=h^{\rm GW}_{+}\epsilon^{+}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{+}}+h^{\rm GW}_{\times}\epsilon^{\times}_{ij}(\boldsymbol{p},\psi)e^{i\phi_{\times}}\;,$ (10)
where $\epsilon^{+}_{ij}$ and $\epsilon^{\times}_{ij}$ are two polarization tensors depending on $\boldsymbol{p}$ and the GW polarization angle $\psi$ (Arzoumanian et al. (2023)), $h_{+}^{\rm GW}$ and $h_{\times}^{\rm GW}$ are their amplitudes, and $\phi_{+}$ and $\phi_{\times}$ are two corresponding phases. We consider the monochromatic GW radiated from a binary system with a circular orbit, the amplitudes $h_{+}^{\rm GW}$ and $h_{\times}^{\rm GW}$ of which can be respectively expressed as
$h_{+}^{\rm GW}=h^{\rm GW}_{0}\frac{1+\cos^{2}\iota}{2}\;,\quad h_{\times}^{\rm GW}=h^{\rm GW}_{0}\cos\iota\;,$ (11)
where $h^{\rm GW}_{0}$ is nearly a constant depending on the chirp mass, the angular orbital frequency and the luminosity distance to the source, and $\iota$ is the inclination angle of the binary. Finally, for the circular orbit, the relation $\varphi_{+}-\varphi_{\times}=\pi/2$ should be satisfied.
We omitted the “star term” in Eq. (8). According to Moore et al. (2017), the ratio between the “star term” and the “Earth term” equals the ratio between the wavelength of the GW and the distance to the source. For a GW frequency of $>10^{-10}\,{\rm Hz}$, the wavelength is $<100\,{\rm pc}$, making the “star term” significantly smaller than the “Earth term” assuming that the star is located at $1$ kpc away from us. Moreover, the “star terms” from different sources are uncorrelated, resembling a kind of random noise, causing the “star terms” being further suppressed after data compression, compared to the “Earth term” that is consistently coherent.
3.2 DPDM with purely gravitational interactions
Analogous to GWs, dark photon fields can also induce metric fluctuations via gravitational effects (Nelson & Scholtz (2011); Khmelnitsky & Rubakov (2014); Nomura et al. (2020)). In this subsection, we derive the angular deflection of light emitted from the star as it passes through DPDM with purely gravitational interactions due to metric perturbations. We should mention that DPDM with a mass below $10^{-22}$ eV is subject to severe constraints from, e.g., the integrated Sachs-Wolfe effect on CMB anisotropies ($\gtrsim 10^{-24}~{}{\rm eV}$) (Hlozek et al. (2015)), the stellar kinematics of dwarf spheroidal galaxies (a few times of $10^{-23}$ eV) (González-Morales et al. (2017); Kendall & Easther (2020)) and the Lyman-$\alpha$ forest ($\gtrsim 10^{-21}~{}{\rm eV}$) (Hui et al. (2017); Armengaud et al. (2017)). The lower bound on ultralight dark matter mass is still under debate, here we treat the local dark matter density $\rho_{\rm DM}$ as a free parameter.
@@ 3.2 DPDM with purely gravitational interactions
where the polarization vector $\boldsymbol{q}(\boldsymbol{x})=(\cos\theta_{q}\cos\varphi_{q}e^{i\phi_{1}}$, $\cos\theta_{q}\sin\varphi_{q}e^{i\phi_{2}},\sin\theta_{q}e^{i\phi_{3}})^{\rm T}$ is parameterized by two polarization angles $\theta_{q}$ and $\varphi_{q}$ and three phases $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, all of which are in general certain functions of $\boldsymbol{x}$. Under the synchronous gauge, the metric perturbation has the following form
$h_{ij}^{\rm DP}(t,\boldsymbol{x})\simeq{\rm Re}\left[\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})e^{2im_{A}t}\right],$ (13)
@@
$\mathcal{H}_{ij}^{\rm DP}(\boldsymbol{x})=-h^{\rm DP}_{0}\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right].$ (14)
The signal amplitude can be written as $h^{\rm DP}_{0}=4\pi GA^{2}_{0}(\boldsymbol{x})$, with $A_{0}(\boldsymbol{x})$ being the amplitude of dark photon field. The amplitude of DPDM is
@@
$h^{\rm DP}_{0}=5.2\times 10^{-15}\left(\frac{\rho_{\rm DM}}{0.4\,{\rm GeV}\,{\rm cm}^{-3}}\right)\left(\frac{10^{-23}\,{\rm eV}}{m_{A}}\right)^{2}.$ (16)
It is known that the metric perturbation $h_{ij}^{\rm DP}$ given in Eq. (13) yields a delay of the arrival time of the pulsar light, which can be probed by PTAs (Nomura et al. (2020); Xue et al. (2022); Wu et al. (2022); Xia et al. (2023)). Here we show that the same metric perturbation could also lead to the deflection of star’s proper position, namely,
$\delta n^{i}_{\rm DP}\simeq-\frac{(\delta^{ik}-n^{i}n^{k})}{2}n^{j}h_{jk}^{\rm DP}({\rm E})+\mathcal{O}(v_{\rm vir})\;,$ (17)
where we have omitted terms relying on the virial velocity $v_{\rm vir}$. The complete gauge-invariant expressions of $\delta n^{i}_{\rm DP}$ can be found in appendix B. The perturbation $h_{jk}^{\rm DP}({\rm E})$ is again calculated at the Earth and we ignored all the “star terms”. In Secs. B.1, B.2 and B.3, we demonstrate that this approximation is valid as long as $m_{A}\geq 10^{-24}$  eV.
3.3 DPDM with additional gauge couplings
DPDM can also interact with ordinary matter if it is associated with ${\rm U(1)}_{B}$ or ${\rm U(1)}_{B-L}$ gauge symmetry. As a consequence, any baryonic matter could experience an oscillating force due to the DPDM background, leading to an acceleration $\boldsymbol{a}$ given by (Pierce et al. (2018))
@@ 3.3 DPDM with additional gauge couplings
which can cause aberrations of the stellar proper motion due to the moving observer (Guo et al. (2019)). The resulting apparent angular deflection can be expressed as
$\delta\boldsymbol{n}_{B/B-L}(t,\boldsymbol{n})=\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})-(\delta\boldsymbol{v}(t,\boldsymbol{x}_{\rm E})\cdot\boldsymbol{n})\boldsymbol{n}\;.$ (20)
where $\boldsymbol{x}_{\rm E}$ is the coordinate of the observer, $\boldsymbol{n}$ is the sky location of the star. Moreover, the dark photon field can also directly lead to stellar motion, with an amplitude approximated by $\delta\boldsymbol{n}^{s}_{B/B-L}\lesssim|\delta\boldsymbol{v}|\cdot t/D$ (with $D$ being the distance between the star and the Earth). However, for the typical observation time $(\sim 1~{}{\rm year})$ and distance $D\gtrsim 1~{}{\rm kpc}$, $\delta\boldsymbol{n}^{s}_{B/B-L}\sim 3\times 10^{-4}\delta{\bm{n}}_{B/B-L}$ is rather tiny and will not be taken into consideration (Guo et al. (2019)).
4 Sensitivity of GW and DPDM searches with Gaia
In this section, we estimate the sensitivity of Gaia-like astrometry for GW and DPDM induced signals. We consider both Gaia and its next-generation upgrade (“XG-Gaia”), where we anticipate that the XG-Gaia will achieve $\mu$as angular resolution while maintaining the same number of observed stars as Gaia does (Vallenari (2018)). The operational parameters of these missions are set as follows:
@@ 4 Sensitivity of GW and DPDM searches with Gaia
where $\sigma_{\rm ss}$ is the noise from a single exposure, [^2] The uncertainties on the proper motion parameters depend on stellar magnitudes $G$. It is expected the final uncertainty $\sigma_{\rm ss}\simeq 100~{}\mu{\rm as}$ for stars with $G\simeq 20$ in Gaia DR5 (Vallenari et al. (2023)). As these stars constitute the vast majority of the Gaia dataset, we will use $\sigma_{\rm ss}=100~{}\mu{\rm as}$ for all stars later on. $\Delta t$ is the observational cadence, $T_{\rm obs}$ is the total observation time, and $N$ is the total number of stars considered in the analysis.
@@
As previously mentioned, we compress the data of $10^{5}$ stars in the same sky area by randomly generating $\widetilde{N}=10^{4}$ virtual stars uniformly distributed across the sky. Consequently, the instrument error for the measurement on each virtual star is rescaled to $\widetilde{\sigma}=\sigma_{\rm ss}/\sqrt{10^{9}/10^{4}}$. To estimate the sensitivity in detecting GWs and DPDM with Gaia-like experiments, we construct a mock dataset by injecting a specific GW or DPDM signal into the virtual stars. In the end of Sec. 3.1, Sec. 3.2 and Sec. 3.3, we demonstrated that the signals we considered are consistently dominated by the “Earth term” for $f_{\rm GW}>10^{-10}$ Hz or $m_{A}>10^{-24}$ eV. Therefore, the astrometric effects on the stars in one area, caused by GWs or DPDM, can be approximated to the global effect on the virtual stars within that area.
@@
In addition to the signal component, in Sec. 2 we have established that the background motion due to stellar proper motion and the acceleration of the SSB is effectively removed by linear subtraction. The constant acceleration resulting from the galaxy’s mass distribution is sufficiently small, thus can be safely ignored. Since linear subtraction is consistently applied to the astrometric data of virtual stars, background motion is not injected into the analysis. To address the time-dependent noise caused by stellar binaries, we adopt two different strategies for Gaia and XG-Gaia:
@@
Figure 2: The 95% exclusion limits from Gaia (green lines) and XG-Gaia (red lines) on the GW strain $h_{0}^{\rm GW}$ from a binary system, the energy density $\rho_{\rm DM}$ of the DPDM with purely gravitational interactions, and the coupling constant $\epsilon$ of $U(1)_{B}/U(1)_{B-L}$ DPDM, respectively. Model parameters regarding all the angles have been integrated out. Solid lines represent the bounds with linear subtraction, while dashed lines correspond to quadratic subtraction. We also use blue lines to denote the results including the time-dependent noise from unresolved binary systems. For comparison, constraints from NANOGrav 15-year (Agazie et al. (2023b); Afzal et al. (2023)), EPTA DR2 (Smarra et al. (2023)), and PPTA DR2 (Xue et al. (2022); Wu et al. (2022)) results are labeled by orange, cyan and purple curves, respectively, with the solid (dashed) curves being the (un)correlated constraints. In the top-right panel, the black dashed line shows the amount of local dark matter density near the Earth as a reference. In the bottom two panels, the horizontal solid lines are the limits from the MICROSCOPE WEP experiment (Touboul et al. (2017, 2022); Amaral et al. (2024); Fayet (2018, 2019)). In addition, the vertical dashed line in each plot refers to the reference frequency $f=1/T_{\rm obs}$ with $T_{\rm obs}=10~{}{\rm year}$ being Gaia’s observation period.
@@
  • For Gaia, we assume that all stellar binaries are unresolvable. Thus we use the PSD $S_{\rm bin}(f)$ calculated in Sec. 2, as shown in Fig. 1.
@@
$\begin{split}\text{(A)}:&\quad\{f_{\rm GW},\;h_{0}^{\rm GW},\;\theta_{p}\;,\varphi_{p},\;\iota,\;\psi,\;\phi_{+}\}\;;\\ \text{(B)}:&\quad\{m_{A},\;\hat{\rho}_{\rm DM},\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\}\;;\\ \text{(C)}:&\quad\left\{m_{A},\;\epsilon^{2}\hat{\rho}_{\rm DM},\;\theta_{q},\;\varphi_{q},\;\phi_{1}\;,\phi_{2}\;,\phi_{3}\right\}\;,\end{split}$
where
$\hat{\rho}_{\rm DM}\equiv\frac{\rho_{\rm DM}}{0.4\,{\rm GeV}\,{\rm cm}^{-3}}.$ (23)
@@
We fix the values of $f_{\rm GW}$ and $m_{A}$, and scan the other model parameters in their individual parameter space in each case. The 95% exclusion limits on the $h_{0}^{\rm GW}$, $\rho_{\rm DM}$ and $\epsilon$ after integrating out all angular parameters are shown in Fig. 2. Some remarks are in order. Firstly, the frequency $f=2/T_{\rm obs}$ approximately marks where the sensitivity curves show significant changes. In the high-frequency range ($f\gtrsim 2/T_{\rm obs}$), the limits obtained using two different subtraction methods are indistinguishable, as expected, whereas linear subtraction offers better sensitivity in the low-frequency range compared to quadratic subtraction. In particular, the 95% exclusion limits are enhanced by more than one order of magnitude at $f_{\rm GW}\lesssim 6\times 10^{-10}\,{\rm Hz}$ for GWs, or at $m_{A}\lesssim 3\times 10^{-24}\,{\rm eV}$ for DPDM. Secondly, the impact of time-dependent noise from unresolved binary systems significantly decreases the sensitivity by up to an order of magnitude. Hence the inclusion of such noise is important in our analysis. In addition, this decrease is more severe in the linear subtraction scheme than in the quadratic subtraction, as we explained in Sec. 2. Lastly, XG-Gaia will possess extended observation duration and improved precision. It will also have enhanced resolution capabilities for binary systems, which is crucial for suppressing the noise from unresolved binaries. As a result, the sensitivity of XG-Gaia to detecting GWs and DPDM is expected to increase by more than three orders of magnitude.
5 Conclusion
Astrometric measurements offer a new approach to detecting signals such as GWs and DPDM in the nHz frequency range. Typically, periodic signals with frequencies below $1/T_{\rm obs}$ are easily mistaken for background noise or systematic uncertainties. Known background sources include (1) the proper motion of stars, (2) stellar acceleration due to the Milky Way’s mass distribution, (3) acceleration from companion stars in binary systems, and (4) the Earth’s acceleration. In Sec. 2, we demonstrate that a linear subtraction method is sufficient to remove all these effects. Compared to the commonly used quadratic subtraction, linear subtraction significantly enhances detection sensitivity in the low-frequency range.
However, other factors, such as telescope movement, changes in satellite orientation, and unknown systematics, are not investigated here. We project detection prospects in this study, assuming all these systematics are properly subtracted. We emphasize that future detection claims will require more detailed studies of these systematics. Furthermore, this work does not explore the angular patterns induced by gravitational waves and dark photon dark matter, which may be crucial for distinguishing between systematics and genuine signals.